OpenStudy (anonymous):

Which of the following is the equation of a line that has a slope of zero and passes through the point (4,6)? x = 4 x = -4 y = -6 y = 6 y = 0

jimthompson5910 (jim_thompson5910):

y = mx+b 6 = (0)(4)+b 6 = 0+b 6 = b b = 6 So the equation of the line is y = 0x+6 which can be simplified to y = 6 So the answer is choice D